State the Fubini’s Theorem for the Triple Integrals and list all six possible order of


integration

1
Expert's answer
2022-01-20T15:34:02-0500

"\text{Fubini's theorem for triple integral states that if f(x,y,z) is continuous on a rectangular}\\
\text{box D}=[a,b]\times[c,d]\times[e,f], \text{then}\\
\int\int\int_D\text{f(x,y,z) dV}=\int^f_e \int^d_c \int^b_a\text{f(x,y,z) dxdydz}.\\
\text{Furthermore, for a,b,c,d,e and f real numbers, the iterated triple integral can be}\\
\text{expressed in six different orderings:}\\
\int^f_e\int^d_c\int^b_a \text{f(x,y,z) dxdydz}=\int^f_e(\int^d_c(\int^b_a \text{f(x,y,z) dx)dy)dz}\\
=\int^d_c(\int^f_e(\int^b_a \text{f(x,y,z) dx)dz)dy}\\
=\int^b_a(\int^f_e(\int^d_c \text{f(x,y,z) dy)dz)dx}\\
=\int^f_e(\int^b_a(\int^d_c \text{f(x,y,z) dy)dx)dz}\\
=\int^d_c(\int^b_a(\int^f_e \text{f(x,y,z) dz)dx)dy}\\
=\int^b_a(\int^d_c(\int^f_e \text{f(x,y,z) dz)dy)dx}\\"